# Strategic resource allocation in memory encoding: An efficiency principle shaping language processing

### v1

@@ Abstract
How is the limited capacity of working memory efficiently used to support human linguistic behaviors? In this paper, we investigate strategic resource allocation as an efficiency principle for memory encoding in sentence processing. The idea is that working memory resources are dynamically and strategically allocated to prioritize novel and unexpected information, enhancing their representations to make them less susceptible to memory decay and interference. Theoretically, from a resource-rational perspective, we argue that this efficiency principle naturally arises from two functional assumptions about working memory, namely, its limited capacity and its noisy representation. Empirically, through naturalistic corpus data, we find converging evidence for strategic resource allocation in the context of dependency locality from both the production and the comprehension side, where non-local dependencies with less predictable antecedents are associated with reduced locality effect. However, our results also reveal considerable cross-linguistic variability, highlighting the need for a closer examination of how strategic resource allocation, as a universal efficiency principle, interacts with language-specific phrase structures.
Keywords: strategic resource allocation; working memory efficiency; dependency locality; cross-linguistic; sentence production; sentence comprehension; naturalistic data
@@ 1 Introduction
Language processing in humans relies on working memory, a cognitive module known for its limited capacity to retain information (Baddeley, 1992; Just & Carpenter, 1992; Fedorenko et al., 2013). A perceived linguistic signal, under this limitation, is at the risk of being lost, as it is rapidly overwhelmed by the continual torrent of new input (Christiansen and Chater, 2016). Meanwhile, language use seems effortless, with sophisticated linguistic representations being dynamically encoded and decoded within milliseconds. This dual nature of working memory raises the question: how is the limited capacity of working memory efficiently used to support human linguistic behaviors?
@@
In this paper, we argue for strategic resource allocation as an efficiency principle for memory encoding in sentence processing: working memory resources are dynamically and strategically allocated to prioritize novel and unexpected information. We argue that this efficiency principle, as a resource-rational theory (Lieder & Griffiths, 2020; Lewis et al., 2014; Gershman et al., 2015), naturally arises from two functional assumptions about working memory: its capacity is limited, and its representations are noisy. To examine this efficiency principle, we report three studies using naturalistic corpus data, where we demonstrate an interaction between strategic resource allocation and the locality effect in processing non-local syntactic dependencies.
2 Background
@@
We first present the theoretical justification and empirical support for our proposal of strategic resource allocation, drawing from the literature on sentence processing and visual working memory. We then review some critical findings about the dependency locality effect, an empirical domain through the lens of which we examined our proposal.
2.1 Strategic Resource Allocation
2.1.1 Theoretical Justification
@@
We argue that working memory resources are strategically allocated in a way that prioritizes novel and unexpected information given the context, an efficiency principle that we refer to as strategic resource allocation. This principle is in line with the resource-rational analysis of human mind (Lieder & Griffiths, 2020; Lewis et al., 2014; Gershman et al., 2015). Grounded in the bounded-rational approach to cognition (Anderson, 1990; Simon, 1955), resource-rational analysis aims to integrate the functional goals of a computational problem into the structural constraints of the underpinning cognitive architectures, providing a linkage between the computational-level and the algorithmic-level theories (Marr, 1982). In other words, instead of looking for an unbounded optimization, resource-rational analysis seeks to explain human behaviors under bounded rationality, that is, to identify an optimal solution that strikes the balance between maximizing the functional utility and adhering to the structural constraints of the cognitive system.
@@
The resource-rational explanation for strategic resource allocation is rooted in two widely held functional assumptions about working memory. First, the capacity of working memory is assumed to be limited. Although the exact nature of this limitation is still under debate, recent models in some non-linguistic domains have shifted from a discrete slot representation (Miller, 1956; Cowan, 2001; Luck & Vogel, 1997; Pashler, 1988) towards a continuous resource-based representation, where the limited capacity can be flexibly allocated across the encoded information (Ma et al., 2014; van den Berg et al., 2012; Brady et al., 2016; Brady and Tenenbaum, 2013; van den Berg & Ma, 2018; Sims et al., 2012; Sims, 2016; Jakob & Gershman, 2023; Bates and Jacobs, 2020).
Figure 1: Conceptual framework of strategic resource allocation in memory encoding. Different linguistic contexts yield different prior distributions for the input $w^{*}$, one being predictable and the other unpredictable. When encoding $w^{*}$, allocating more memory resources results in sharper encoding with less uncertainty, such that more probability mass is concentrated around the true input. Upon retrieval, although high surprisal $w^{*}$ is less likely to be successfully reconstructed, the one encoded with higher resources is reconstructed with less error than the one encoded with lower resources.
@@
Second, memory representation is full of noise and uncertainties, with unpredictable corruption of the veridical forms of sensory input, resulting in distorted representations that undermine behavioral performance such as inaccurate recall and comprehension (Ma et al., 2014; Gibson et al., 2013; Levy, 2008b; Brady et al., 2024). Given the noisy input, the true state of a past input is assumed to be inaccessible, and memory recall is often implemented as an inferential process from uncertainties using the statistical structure of long-term knowledge, a process often mathematically characterized by Bayes’ rule (Futrell, Gibson, and Levy, 2020; Gibson et al., 2013; Levy, 2008b; Ryskin et al., 2021; Bays et al., 2024):
$p(w\mid r)\propto p_{M}(r\mid w)p(w).$ (1)
Specifically, as shown in Figure 1, a past input unit $w^{*}$ is first encoded as a distribution over representations $r$ over all words with uncertainty. Later in retrieval, the input unit is reconstructed from this noisy memory representation $r$ using prior knowledge $p(w)$. In this process, units that are more probable in the prior are more likely to be accurately reconstructed, resulting in higher posterior recall accuracy.
The two assumptions above then naturally give rise to the following optimization problem: how to maximize memory accuracy under the constraint of limited resources? We argue that strategic resource allocation can be seen as a rational solution to this optimization problem given the reconstruction mechanism described above. Importantly, we assume that the amount of memory resources allocated to a unit is proportional to the precision of the encoded distribution. That is, as illustrated in Figure 1, more resources allocated to encode $w^{*}$ results in a sharper distribution with less uncertainty, such that more probability mass is concentrated around the true input $w^{*}$.
@@
Now, consider a hypothetical situation in which the memory resources are uniformly distributed across all linguistic units regardless of their surprisal. Under this uniform distribution, each unit will receive an identical encoding distribution. However, due to the differences in reconstructability with regard to prior surprisal, the same encoding distribution would yield different retrieval accuracy, disproportionally exerting impact on high surprisal units, reducing their posterior accuracy more significantly than low surprisal units. Therefore, a uniform distribution of memory resources is not the most efficient way to go, and leaves substantial room for improving the overall posterior accuracy.
@@
An alternative approach, which is the one we explore in the current study, is to allocate resources strategically based on a unit’s predictability from context. As mentioned above, we propose that high surprisal linguistic units a priori are prioritized in working memory, receiving more resources than low surprisal ones. That means the prioritized, more surprising units will receive sharper encoding, resulting in less erroneous retrieval. This asymmetric allocation of memory resources is more efficient than the uniform strategy described earlier, in the sense that it achieves higher accuracy across inputs. By sacrificing a slight reduction in posterior retrieval accuracy for low surprisal units, significant gains can be achieved for high surprisal ones by preventing these irrecontructable units from being distorted in the first place. Put simply, when only a limited number of linguistic units can be encoded with minimal distortion, it is more important to encode the more surprising and less reconstructable ones.
@@
Beyond memory, strategic resource allocation aligns with theories such as predictive coding, free energy principle, and implicit learning. At the neural level, the predictive coding mechanism (Murray et al., 2002; Rao & Ballard, 1999; Aitchison and Lengyel, 2017; Blank and Davis, 2016; Sohoglu & Davis, 2020, 2016; Gagnepain et al., 2012) and the free energy principle (Friston, 2005, 2010; Gershman, 2019) hold that the brain seeks to minimize its prediction error, or surprise, as a way to optimize its internal model of the external environment. This principle is implemented by encoding prediction errors rather than the raw sensory input in neural signals. At the behavioral level, implicit learning theories often hold that learning is error-driven, with considerable empirical support that larger prediction errors lead to greater learning effect (Chang et al., 2006; Scheepers, 2003; Hartsuiker & Kolk, 1998; Bock, 1986; Ferreira, 2003; Jaeger & Snider, 2013; Xu & Futrell, 2024; Rumelhart et al., 1986; Elman, 1990; Courville et al., 2006; Wagner & Rescorla, 1972). Taken together, all these theories delivered a similar implication for our proposal in the domain of working memory. That is, when predictions conflict with the actual perceptual input (that is, when there is high surprisal input), it signals the need for comprehenders to update their mental model in order to make more accurate predictions in the future. Given this critical role of more surprising linguistic units in refining the mental model, more cognitive resources should be allocated to them.
2.1.2 Empirical Support

### v2

@@ Abstract
How is the limited capacity of working memory efficiently used to support human linguistic behaviors? In this paper, we propose Strategic Resource Allocation (SRA) as an efficiency principle for memory encoding in sentence processing. The idea is that working memory resources are dynamically and strategically allocated to prioritize novel and unexpected information. From a resource-rational perspective, we argue that SRA is the principled solution to a computational problem posed by two functional assumptions about working memory, namely its limited capacity and its noisy representation. Specifically, working memory needs to minimize the retrieval error of past inputs under the constraint of limited memory resources, an optimization problem whose solution is to allocate more resources to encode more surprising inputs with higher precision. One of the critical consequences of SRA is that surprising inputs are encoded with enhanced representations, and therefore are less susceptible to memory decay and interference. Empirically, through naturalistic corpus data, we find converging evidence for SRA in the context of dependency locality from both production and comprehension, where non-local dependencies with less predictable antecedents are associated with reduced locality effect. However, our results also reveal considerable cross-linguistic variability, suggesting the need for a closer examination of how SRA, as a domain-general memory efficiency principle, interacts with language-specific phrase structures. SRA highlights the critical role of representational uncertainty in understanding memory encoding. It also reimages the effects of surprisal and entropy on processing difficulty from the perspective of efficient memory encoding.
Keywords: strategic resource allocation; working memory efficiency; resource-rational; dependency locality; cross-linguistic; sentence processing; naturalistic data
@@ 1 Introduction
Language processing in humans relies on working memory, a cognitive module known for its limited capacity to retain information (Baddeley, 1992; Just & Carpenter, 1992; Fedorenko et al., 2013). Under this limitation, a linguistic signal, once perceived, is at the risk of being lost, rapidly overwhelmed by the continual torrent of new inputs (Christiansen and Chater, 2016). Meanwhile, language use seems effortless, with sophisticated linguistic representations being dynamically encoded and decoded within milliseconds. This dual nature of working memory raises the question: how is the limited capacity of working memory efficiently used to support human linguistic behaviors?
@@
In this paper, we propose Strategic Resource Allocation (SRA) as an efficiency principle for memory encoding in sentence processing. Specifically, working memory resources are dynamically and strategically allocated to prioritize novel and unexpected information. We argue that this efficiency principle, as a resource-rational theory (Lieder & Griffiths, 2020; Lewis et al., 2014; Gershman et al., 2015), naturally arises as the solution to a computational problem posed by two functional assumptions about working memory: its capacity is limited, and its representations are noisy. To examine this efficiency principle, we report three studies using naturalistic corpus data, where we demonstrate empirical support for strategic resource allocation through the lens of the locality effect in processing non-local syntactic dependencies.
2 Strategic Resource Allocation (SRA)
@@
We first present the theoretical justification and existing empirical evidence for our proposal of Strategic Resource Allocation (SRA), drawing from the literature on sentence processing and psychophysics.
2.1 Theoretical Proposal
We propose that working memory resources are strategically allocated in a way that prioritizes novel and unexpected information given the context, an efficiency principle that we refer to as Strategic Resource Allocation (SRA):
  1. (1)
    Strategic Resource Allocation (SRA) in memory encoding:
    Principle. Working memory resources are dynamically and strategically allocated in a way that prioritizes linguistic units that are unexpected and surprising given the context.
    Core Prediction. The encoding of more surprising units is enhanced, resulting in more robust memory representations that are less susceptible to memory interference or decay.
@@ 2.1 Theoretical Proposal
This principle is in line with the resource-rational analysis of human mind (Lieder & Griffiths, 2020; Lewis et al., 2014; Gershman et al., 2015). Grounded in the bounded-rational approach to cognition (Anderson, 1990; Simon, 1955), resource-rational analysis aims to integrate the functional goals of a computational problem into the structural constraints of the underpinning cognitive architecture, providing a linkage between the computational-level and the algorithmic-level theories (Marr, 1982). In other words, instead of looking for an unbounded optimization, resource-rational analysis seeks to explain human behaviors under bounded rationality, that is, to identify an optimal solution that strikes the balance between maximizing the functional utility and adhering to the structural constraints of the cognitive system.
In this section, we will first outline the computational problem faced by working memory: to infer past information from uncertainty with maximal accuracy under the constraint of limited memory resources. We will then explain how SRA provides an optimal solution to this computational problem.
2.1.1 Inferring from uncertainty: A computational problem of working memory
Figure 1: Working memory processes under the probabilistic framework. The linguistic input $w$ is encoded with noisy internal memory representation $r$. Higher memory resources results in sharper representation concentrated around the true input with less uncertainty. Memory retrieval can be considered an inference process reconstructing linguistic input based on noisy representation.
@@ 2.1.1 Inferring from uncertainty: A computational problem of working memory
As already mentioned, the resource-rational explanation for SRA is rooted in a computation problem posed by two functional assumptions about working memory. First, the capacity of working memory is limited. Although the exact nature of this limitation is still under debate, recent models in some non-linguistic domains have shifted from a discrete slot representation (Miller, 1956; Cowan, 2001; Luck & Vogel, 1997; Pashler, 1988) towards a continuous resource-based representation, where the limited resources can be flexibly allocated across the encoded information (Ma et al., 2014; van den Berg et al., 2012; Brady et al., 2016; Brady and Tenenbaum, 2013; van den Berg & Ma, 2018; Sims et al., 2012; Sims, 2016; Jakob & Gershman, 2023; Bates and Jacobs, 2020).
Figure 2: Conceptual framework of strategic resource allocation (SRA) versus uniform resource allocation in memory encoding. Compared to uniform allocation, SRA holds that more surprising units receive more memory resources, therefore less representational uncertainty in its encoding distribution. Upon retrieval, more surprising units are less reconstructable based on prior compared to less surprising ones. But SRA leads to lower retrieval error overall by improving the reconstruction of high-surprisal units with minimal loss to the retrieval accuracy on low-surprisal units.
@@
Second, memory representation is full of noise and uncertainty, with unpredictable corruption in the veridical forms of sensory input, resulting in distorted representations that undermine behavioral performance such as inaccurate recall and illusive comprehension (Ma et al., 2014; Gibson et al., 2013; Levy, 2008b; Brady et al., 2024; F. Ferreira et al., 2002). This uncertainty is often represented under the probabilistic framework. As shown in Figure 1, when a linguistic input $w$ is received, it can be encoded into an internal representation through certain memory model $r=M(w)$. This $r$ is a probabilistic distribution centered around the true value of that input, such that the true input bears the highest probability in the encoding distribution compared to other alternatives. [^1] In many psychophysics studies, the encoded representation is assumed to be a specific stimulus value that is generated from certain probabilistic distribution. Here in our work, we take a different assumption and postulate that what has been encoded, instead of a specific stimulus value, is the distribution itself, either through sampling (Hoover et al., 2023) or through probabilistic population codes (Ma et al., 2006). Given this noisy representation, the true state of a past input is inaccessible, and memory recall, decoding, or retrieval, is effectively an inferential process that reconstructs past input from uncertainty using the statistical structure of long-term knowledge. The results of this process are often mathematically characterized using Bayes’ rule (e.g., Futrell, Gibson, and Levy, 2020; Gibson et al., 2013; Levy, 2008b; Ryskin et al., 2021; Bays et al., 2024):
$p(\hat{w}\mid r)\propto p_{M}(r\mid w)p(w).$ (1)
The equation describes a rational Bayesian decoder which infers the input from a specific memory representation $r$ integrating prior knowledge $p(w)$, yielding a posterior distribution $p(\hat{w}\mid r)$. Then, marginalizing over all possible values of $r$, the distribution on the reconstructed word given the true input $w^{*}$ is
$p(\hat{w}\mid w^{*})=\int p(\hat{w}\mid r)p(r\mid w^{*})dr.$ (2)
In this inferential process, inputs that are more probable in the prior are more likely to be accurately reconstructed, resulting in higher retrieval accuracy. See Appendix A.1 for detailed mathematical formalization of this probabilistic memory encoding and retrieval process.
The two assumptions above naturally give rise to the following optimization challenge: how to maximize memory accuracy under the constraint of limited resources? At the core of this challenge lies an efficiency problem for two reasons. First, there is a functional goal, which is memory accuracy, against which the working memory performance is evaluated. Such a functional nature situates the current proposal under the rationalist approach to human mind. Second, working memory has internal constraints, in the sense that there is something it cannot achieve due to the cost from its own structure. Without such constraints, there would be no reason to look for an efficient implementation of a functional goal. The acknowledgment of system-internal cost, therefore, further situates the current proposal under resource-rationality. [^2] In this paper, we simply represent the internal cost of working memory as a computational bound (i.e., the total amount of available memory resources) within which a given task such as memory retrieval is imperfectly optimized, an approach that has been termed bounded optimality Icard (2023). Next, we will explain how SRA provides a principled solution to the computational challenge outlined above.
2.1.2 SRA as a resource-rational solution
First of all, an important assumption we make is that the precision of the encoded distribution is proportional to the amount of memory resources allocated to an input unit. That is, as illustrated in Figure 1, more resources allocated to encode $w$ results in sharper distribution with less uncertainty, such that more probability mass is concentrated around the true input value $w$. Despite its lack of attention in the field of sentence processing, this assumption has been widely entertained in the literature of psychophysics (e.g., Bates and Jacobs, 2020; Bays et al., 2024, 2009; Ma et al., 2014). [^3] One of the biggest challenges to apply such an encoding distribution in the domain of language is the specification of hypothesis space. In psychophysics, the hypothesis space is usually a quantitative spectrum that can be objectively specified based on certain physical features. But for language, the linguistic inputs are discrete units. Nowadays, with the advance of modern NLP techniques, this challenge has been significantly mitigated given the distributive word representations such as word embeddings. However, it is still nontrivial work to figure out what kind of probabilistic distribution should be applied to word embedding space. See Appendix A.1 for a Gaussian approximation to the probabilistic memory processes. As shown below, such a relationship between allocated resources and representational uncertainty lays the foundation for the derivation of SRA.
@@ 2.1.2 SRA as a resource-rational solution
Before demonstrating the rationale behind SRA, let us first consider a naive strategy in which memory resources are uniformly distributed across all linguistic units regardless of the statistical structure in the context (Figure 2, bottom panel). Under this uniform distribution, each input unit will receive an encoding distribution with identical precision. However, under the influence of prior, inputs that are more surprising under the prior will be less reconstructable, and the retrieval distribution will be more drawn towards the prior. Due to this difference in reconstructability, the same encoding distribution would yield different retrieval accuracy, disproportionally exerting impact on high surprisal inputs, reducing their retrieval accuracy more significantly than low surprisal ones. Therefore, a uniform distribution of memory resources is not the most efficient way to go for memory encoding, leaving substantial room for the improvement of overall retrieval accuracy.
@@
Now, consider SRA, the strategic allocation of memory resources (Figure 2, top panel). Recall that the idea is to strategically allocate more resources on linguistic units of higher surprisal a priori. That means, the prioritized more surprising units will receive sharper encoding with higher precision. This asymmetric allocation of resources is more efficient than the uniform strategy described in the last paragraph, since it achieves higher accuracy on average across inputs. By sacrificing a slight reduction in retrieval accuracy for low surprisal units, significant gains can be achieved for high surprisal ones by preventing these irrecontructable units from being distorted in the first place. Put simply, when only a limited number of linguistic units can be encoded with minimal distortion, it is more important to encode the more surprising and less reconstructable ones (See Appendix A.2 for mathematical derivation). [^4] By having an encoding strategy that optimizes faithful reconstruction of the true input, an implicit assumption we made is that the input signal is considered error-free. If the input itself contains errors (e.g., when there are speech errors produced by the speaker), a reconstruction that is strongly influenced by the prior may actually be preferred so that the signal errors can be corrected. How to deal with the errors in input signals is an important online processing task. But in the current proposal, we choose to analyze working memory as a system of information storage, whose main goal is to accurately encode and decode the information it receives (cf. Hasson et al., 2015).
@@
Beyond memory, SRA aligns with theories such as predictive coding, free energy principle, and implicit learning. At the neural level, the predictive coding mechanism (Murray et al., 2002; Rao & Ballard, 1999; Aitchison and Lengyel, 2017; Blank and Davis, 2016; Sohoglu & Davis, 2020, 2016; Gagnepain et al., 2012) and the free energy principle (Friston, 2005, 2010; Gershman, 2019) hold that the brain seeks to minimize its prediction error, or surprise, as a way to optimize its internal model of the external environment. This principle is implemented by encoding prediction errors rather than the raw sensory input in neural signals. At the behavioral level, implicit learning theories often hold that learning is error-driven, with considerable empirical support showing that larger prediction errors lead to greater learning effect (Chang et al., 2006; Scheepers, 2003; Hartsuiker & Kolk, 1998; Bock, 1986; V. S. Ferreira, 2003; Jaeger & Snider, 2013; Xu & Futrell, 2024; Rumelhart et al., 1986; Elman, 1990; Courville et al., 2006; Wagner & Rescorla, 1972). Taken together, all these theories delivered a similar implication for our proposal in the domain of working memory. That is, when predictions conflict with the actual perceptual input (that is, when there is high surprisal input), it signals the need for comprehenders to update their mental model in order to make more accurate predictions in the future. Given this critical role of more surprising linguistic units in refining the mental model, it is reasonable to allocate more memory resources to them.
2.2 Predictability-Precision Trade-Off
SRA can also be construed as a predictability-precision trade-off, as illustrated in Figure 3. Imagine there are multiple linguistic units to be stored in working memory, with the goal of minimizing their total retrieval error (or, maximizing the retrieval accuracy). In most cases, for each unit, higher memory resources suggest higher encoding precision, which lead to lower retrieval error. [^5] In some borderline cases, where the input word is too close to the prior prediction and the prior precision is too unreliable, the retrieval error may not monotonically decrease with increasing encoding precision. However, as shown in Appendix A.2, in either case, the strategic resource allocation should still hold in the sense that more resources should be allocated to high surprisal units in order to minimize the expected total error. However, given a fixed amount of resources, allocating more to one unit necessarily reduces what is available to others. Therefore, the gain of retrieval accuracy for one unit necessarily comes with the loss for others. Consequently, to minimize total retrieval error, the optimal distribution of a fixed amount of resources should balance the gain and the loss in retrieval accuracy across units.
Figure 3: Retrieval error as a function of encoding precision for high-surprisal and low-surprisal inputs. The optimal encoding strategy balances the potential gain and loss in retrieval accuracy across linguistic inputs.
Where, then, does the balance hold? The intuition is illustrated in Figure 3. First, let us start from the uniform distribution of resources, where both high- and low-surprisal inputs receive the same amount of resources, and are thus encoded with the same degree of precisions (i.e., gray dots in Figure 3). Importantly, the retrieval error decreases faster for high-surprisal input. Therefore, at this point of uniform distribution of resources, there is a momentum to redistribute more resources to high-surprisal input. This is because such a redistribution towards high-surprisal input reduces the error faster. In other words, given a fixed amount of resources, the pressure to lower the overall retrieval error will push the encoding precision of high-surprisal input to increase from a uniform distribution of resources, and vice versa for low-surprisal one. This redistribution continues until there is balanced marginal effects, that is, when the slopes of the two error functions are equal, as in Figure 3 (see Appendix A.2 for details).
2.3 Precision, Accuracy, and Robustness
It is important to point out that the key prediction of SRA is not simply about the mean accuracy of memory retrieval. As shown in Figure 2 (top panel), both predictable and unpredictable units can achieve relatively high accuracy with respect to the mean of retrieval distribution. For predictable input, this is supported by prior knowledge; for unpredictable input, this is achieved by more precise encoding representation. In fact, one of the critical consequences of SRA is that the retrieval mean accuracy should remain approximately similar across different linguistic units. [^6] SRA does not necessarily predict the retrieval mean accuracy to be an absolute constant across all linguistic units. In fact, the accuracy for low surprisal units may still be higher than high surprisal ones in Bayesian inference. However, due to SRA, this difference can be reduced compared to a naive encoding strategy such as uniform resource allocation.
But what does differ is the precision, or uncertainty in the representation. By allocating more resources, more surprising linguistic units are encoded with lower uncertainty (see the sharper distributions with high resources in Figure 2). We thus propose that the uncertainty in the memory representation, rather than being linked to retrieval accuracy, is more directly related to memory robustness. That is, memory representation of higher robustness is less susceptible to the interference from other elements in memory. It of course remains a debatable question what the linking hypothesis is for representational uncertainty, but the linkage between uncertainty and robustness gives us a working hypothesis that is readily testable, as shown below in the rest of this paper. We will return to this point later in General Discussion 7.2.
Our theoretical framework of SRA alludes to three effects on memory encoding precision:
  1. (2)
    Three predicted effects of strategic resource allocation on encoding precision
    1. a.
      Effect of input surprisal
      Surprising linguistic units bear higher encoding precision, resulting in more robust memory representation against interference.
    2. b.
      Effect of memory constraint
      More available memory resources result in higher encoding precision overall.
    3. c.
      Effect of prior precision
      Precision of prior prediction does not necessarily increase or decrease encoding precision.
The most important prediction of SRA is (2a), the effect of input surprisal. As outlined earlier, the optimal strategy to minimize overall retrieval error is to allocate more memory resources to encode surprising input. This strategy results in higher precision in the representation, thus higher memory robustness against interference. This is the critical prediction that we are going to examine empirically in the current study.
For the effect of memory constraint (2b), SRA predicts that more available memory resources results in higher encoding precision in general. This will also lead to higher memory robustness and more accurate retrieval overall.
For the precision of prior prediction (2c), its effect on encoding precision is in fact less straightforward. When the true input is very close to the prior prediction, it is indeed possible that less uncertainty in the prior can better support memory retrieval, thus less precise encoding is needed. However, when the true input is far from the prior prediction, it is not necessarily the case that more precise prior can still support better retrieval. We will discuss this effect in more detail and its implication for the effect of prediction entropy on processing difficulty in General Discussion 7.1.
2.4 Some Existing Empirical Evidence